OpenStudy (anonymous):

Help please MEDLE & FAN!:) Directions: Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. Blake has entered a mountain bike competition at Starvation Flats Bike Park. Placement in the competition is determined by a person's average speed over four laps around the course. The course itself is of a mile long. If Blake completed the four laps in 12 minutes, then his average speed around the track is ______ miles per hour.

OpenStudy (anonymous):

@iGreen @dillon754 @PaulaLovesSchool13

OpenStudy (igreen):

The course is 1 mile long. Blake ran around the course 4 times in 12 minutes. That's 4 miles in 12 minutes. 60 / 12 = 5. 4 * 5 = 20 miles. 12 * 5 = 60 minutes = 1 hour. So Brandon runs 20 miles an hour.

OpenStudy (anonymous):

ok thank you!

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

its 8 miles ph

OpenStudy (anonymous):

yes its 8.